In 2019 KRJ International reported earnings before taxes of $100,000 and tax expense of $35,000, leaving net income of $65,000. During the year it sold a warehouse with a net book value of $23,431 for $13,016; the resulting gain or loss on the sale had been included in the calculation of earnings before taxes. Recalculate the company’s net income with the unusual gain or loss removed. PLEASE SHOW STEPS
Answer #1

Net loss = 23,431 - 13,016 = 10,415

Tax rate = (Tax expense / Earnings before tax) * 100

Tax rate = (35,000 / 100,000) * 100

Tax rate = 35%

To remove unusual loss, we need to add back the loss the earnings before tax

Earnings before tax = 100,000 + 10,415 = $110,415

Net income = Earnings before tax(1 - tax)

Net income = 110,415(1 - 0.35)

Net income = $71,769.75